OpenStudy (anonymous):

The following function represents the profit P(n), in dollars, that a concert promoter makes by selling tickets for n dollars each: P(n) = -250n^2 + 2,500n - 5,250 Part A: What are the zeroes of the above function, and what do they represent? Show your work.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

What does n represent?

OpenStudy (anonymous):

dollars

OpenStudy (solomonzelman):

precisely, n, is a price per ticket.

OpenStudy (solomonzelman):

What does P(n) represent?

OpenStudy (anonymous):

profit

OpenStudy (solomonzelman):

Yes, when P(n) is 0, that means that the profit is zero (i.e. made nothing), right?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

The x-intercepts of this function are the values of n (or the value of the price per ticket) with which the P(n) (or the profit) is 0. (There are going to be 2 x-intercepts)

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

So, if you find the x-intercepts, you are going to find the 2 possible values of a price per ticket with which the promoter will NOT profit (but will not lose).

OpenStudy (solomonzelman):

So, do you see what these x-intercepts would mean?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Ok, now you have to solve for these x-intercepts.

OpenStudy (anonymous):

the x intercepts are 3 and 7?

OpenStudy (solomonzelman):

yes, very good!

OpenStudy (solomonzelman):

Perfect!

OpenStudy (solomonzelman):

So, when the ticket costs 3$ or 7$ , then the promoter doesn't gain or lose.

OpenStudy (anonymous):

now what?

OpenStudy (solomonzelman):

when the *average* ticket I should say

OpenStudy (solomonzelman):

Well, all you had to do is to find these zero (these x-intercepts), and to then say what these x-intercepts represent in your situtation.

OpenStudy (solomonzelman):

you found x=3, and x=7 ((( and that is correct ))) now you have to say what do these intercepts mean/do/represent?

OpenStudy (solomonzelman):

just say what we have agreed on in the beginning.

OpenStudy (anonymous):

can you help me with part B?

OpenStudy (solomonzelman):

what is part B?

OpenStudy (solomonzelman):

I can try my best with it:)

OpenStudy (anonymous):

Part B: Find the maximum profit by completing the square of the function P(n). Show the steps of your work.

OpenStudy (solomonzelman):

Oh, good

OpenStudy (solomonzelman):

``` SIDE - NOTE: Most people would say that this question requires calculus, but in a case of a parabola (a quadratic does not). (this what most people say is a mistake). ``` ALL YOU GOT TO DO: You have to write your P(n) in a vertex form, and the vertex is going to be the maximum.

OpenStudy (anonymous):

so (5,1000)?

OpenStudy (solomonzelman):

Let me see.... i wil be typing, and if it is right I will post the work, and if it is not, then I will ask you to redo it and follow along with me.

OpenStudy (solomonzelman):

hope that is fair.

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle P(n) = -250{\rm n}^2 + 2,500{\rm n} - 5,250 }\) \(\large\color{black}{ \displaystyle P(n) = -250({\rm n}^2 -10{\rm n}) - 5,250 }\) \(\large\color{black}{ \displaystyle P(n) = -250({\rm n}^2 -10{\rm n}+25-25) - 5,250 }\) \(\large\color{black}{ \displaystyle P(n) = -250({\rm n}^2 -10{\rm n}+25)+(-250)(-25) - 5,250 }\) \(\large\color{black}{ \displaystyle P(n) = -250({\rm n}^2 -10{\rm n}+25)+1000 }\) \(\large\color{black}{ \displaystyle P(n) = -250({\rm n}-5)^2+1000 }\)

OpenStudy (solomonzelman):

yes, that is correct

OpenStudy (solomonzelman):

And quite quick

OpenStudy (anonymous):

1 graphed it

OpenStudy (solomonzelman):

oh, you graphed it... good, but it is very good (I would insist that even required) to know that without graphing.... in any case though, lets do a little review of what you have actually found just now,

OpenStudy (anonymous):

the axis of symmetry would be 5 right? (Thats part C)

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

yes part C is correct

OpenStudy (solomonzelman):

now the review of what does the vertex of (5,1000) actually mean in this case.

OpenStudy (anonymous):

thats the maximum about of profit the promoter can make?

OpenStudy (solomonzelman):

that is you are refering to the point to a 5 or a 1000 ?

OpenStudy (anonymous):

1000 makes more sense 1 think

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

and what does 5 mean, can you tell me?

OpenStudy (anonymous):

the amount the tickets cost?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

So when the average ticket costs 5$, you get a maximum profit of 1000

OpenStudy (anonymous):

makes sense

OpenStudy (solomonzelman):

A quite small, if not miserable profit in a real life situation... but won't be talking about a real life situation because such situations are never modeled with quadratics in real life....

OpenStudy (solomonzelman):

yeah, so you got any questions regarding any of the PARTS ?

OpenStudy (anonymous):

nope thats it

OpenStudy (anonymous):

as usual you're the most helpful person on the whole site. Thanks!

OpenStudy (solomonzelman):

\(\large{\bbox[5pt, cyan ,border:2px solid black ]{ \rm Thanks! }}\)

OpenStudy (solomonzelman):

I wouldn't say I am most helpful, but won't deny some use.; you are always welcome!